OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 3) and (–2, 6)?

OpenStudy (anonymous):

m= (6-3)/(-2-2) =3/-4 =-3/4

OpenStudy (anonymous):

the slope is -3/4